OpenStudy (anonymous):

Find the area rhombus with a perimeter of 52 inches and a diagonal of 24 inches

OpenStudy (anonymous):

perimeter = 52 so each side = 52/4 = 13 inch side of longest diagonal = 24 a right angled triangle is formed by sides 13, (24/2) and (Other diagonal/2) 5, 12, 13 is triplet so other diagonal half = 5 × 2 = 10 area = product od diagonal length/2 = 24 × 10/2 = 240/2 in² = 120 in²

OpenStudy (anonymous):

hope this helped

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

your welcome